OpenStudy (anonymous):

There are two ions, K+ and Cl-. Find the Zeff for each if the valence electrons contribute 0, and the core electrons contribute 1. Now calculate the Zeff using Slater's rules. Thanks in Advance! First person to write it out in a way I can understand gets the medal!

OpenStudy (anonymous):

valence electrons contribute 0, and core electrons contribute 1 to "S".

OpenStudy (aaronq):

\(Z_{eff}=Z-S\)

OpenStudy (aaronq):

for slater's rules check this: http://intro.chem.okstate.edu/WorkshopFolder/SlaterRule.html

OpenStudy (anonymous):

I understand slater's rules, but I don't understand why the effective nuclear charge will be different for these two ions. They have the exact same electron configuration.

OpenStudy (aaronq):

S = 0.35*x + 0.85*y +z x= valence electrons y= level below valence level z= the rest of the electrons

OpenStudy (anonymous):

I thought x was valence electrons -1

OpenStudy (anonymous):

I always understood slater's rule like this: S= 0.35x + 0.85y +1z x= # of electrons in the group we are calculating the zeff for -1 y= # of electrons in the next group z= # of all the other electrons

OpenStudy (aaronq):

oh sorry, you're right, you're subtracting one from x, so x-1

OpenStudy (anonymous):

Using that, I would calculate the Zeff felt by the valence electrons of Na (ground state) like this: Electron Config: \[1s^2 2s^2 2p^6 3s^1\] \[S= 0(0.35) + 8(0.85) + 2(1)\] \[Z _{eff} = Z - S\] \[Z = 11\] \[S = 8.8\] \[Z _{eff} = 11 - 8.8\] \[Z _{eff} = 2.2+\]

OpenStudy (aaronq):

that looks right, but why are you doing Na not K? i have to go, if you haven't figured it out, i'll comeback

OpenStudy (anonymous):

Alright thanks, I was just testing it out with the ground state of sodium. I need to figure out how to apply those methods for K+ and Cl-.

OpenStudy (aaronq):

just to make sure, you're finding \(Z_{eff}\) for the valence e, right? \(K^+\) \(1s^22s^22p^63s^23p^6\) S= 0.35(0) + 0.85(7) +10 \(Z_{eff}\)=19-(0.35(0) + 0.85(7) +10)=2.2 \(Cl^-\) \(1s^22s^22p^63s^23p^6\) \(Z_{eff}\)=17-(0.35(0) + 0.85(7) +10)=1.05

OpenStudy (aaronq):

sorry, \(Z_{eff}=3.05\;for\;K^+\)

OpenStudy (anonymous):

Why did you multiply 0.35 by 0?

OpenStudy (aaronq):

omg, sorry i just got a new cat and I'm allergic to it, so i've been taking anti-histamines and they make me drowsy. for K+ Zeff=19-(0.35(7) + 0.85(8) +2)= 7.75 for Cl- Zeff=17-(0.35(7) + 0.85(8) +2)= 5.75

OpenStudy (anonymous):

Ah, I understand now! Thank you so much! :)

OpenStudy (anonymous):

Good luck with your new cat :P.

OpenStudy (aaronq):

okay cool! no problem. haha thanks !